# Probabilistic Super-Resolution for Urban Micrometeorology via a Schrödinger Bridge

### v1

@@ 4.2 Variance for ensemble SR inference
Figure 5 shows scatter plots of RMSE versus spread and rank histograms. Both the DM and SM exhibit spread–skill ratios less than 1 and U-shaped rank histograms, indicating ensemble underdispersion. Such underdispersion is common in deep learning models for meteorological problems [e.g., 25], motivating the development of models with larger spread. Compared with the DM, the SM is better calibrated: its point cloud lies closer to the diagonal, with a spread–skill ratio (Spread/RMSE) of 0.658, larger than 0.649 for the DM. Rank histograms are also flatter for the SM; the Jensen–Shannon distance [9] from a perfectly flat uniform distribution is 0.186 for the SM, smaller than 0.238 for the DM.
Figure 5: (Left) Scatter plots of ensemble spread versus ensemble mean RMSE; (Right) rank histograms. (Top) The DM with $N_{T}=50$; (Bottom) the SM with $N_{T}=10$. The number of ensemble members was 64. In the left column, mean spread–skill ratios (Spread/RMSE) are shown in the figure titles, and in the right column, Jensen–Shannon distances [JS distances; 9] from a uniform distribution are shown. The closer the spread–skill ratio is to 1 and the closer the Jensen–Shannon distance is to 0, the more appropriate the ensemble variance.
To assess significance, we used the experimental results from U-Nets initialized with different random weights and calculated the means and standard deviations of the spread–skill ratio and the Jensen–Shannon distance (Table 2). The spread–skill ratio for the SM is closer to 1 than that for the DM, and the Jensen–Shannon distance for the SM is closer to 0. Thus, ensemble inference with the SM yields more appropriate statistics, exhibiting milder underdispersion than the DM.
Table 2: Spread–skill ratios and Jensen–Shannon distances for the DM and SM. These quantities were computed by performing 64-member inference for each ground truth and then averaged over the test data (540 sets). Standard deviations were obtained as in Table 1.
Spread–Skill Ratio (Spread/RMSE) Jensen–Shannon Distance
DM (64 Members, $N_{T}=50$) 0.641 $\pm$ 0.024 0.234 $\pm$ 0.025
SM (64 Members, $N_{T}=10$) 0.657 $\pm$ 0.013 0.189 $\pm$ 0.015
@@ 4.3 Total inference time
Finally, we discuss the total inference time for 60-min predictions. We first report the wall-clock times for physics-based MSSG simulations, measured on the Earth Simulator at the Japan Agency for Marine-Earth Science and Technology (JAMSTEC), equipped with AMD EPYC 7742 CPUs. On average, HR simulations required 206 min with 256 CPU cores, whereas LR simulations required 6.19 min with 40 CPU cores [51]. We then measured the average SM inference time on a local workstation equipped with an NVIDIA L40S GPU. For 60-min LR data (60 sets), single-member inference took 7.29 s. Thus, the hybrid method combining LR simulation and SM inference completed 60-min predictions of 2-m temperature in 6.31 min, reducing HR computation time to 3.06% (a 32.7-fold speedup). This factor is comparable to values reported for recent surrogate models in urban airflow simulations [37, 31].
@@ 5 Conclusions
We applied a Schrödinger-bridge model [SM; 7] to super-resolve urban 2-m temperature and showed that the SM achieves accuracy comparable to that of a diffusion model [DM; 15], while requiring only about one-fifth the computational cost. Using spread–skill ratios and rank histograms, we further showed that SM ensembles yield larger spreads and more appropriate statistics than DM ensembles. While DMs have been actively applied in meteorology [20, 13, 25, 36, 43], our results motivate the exploration of deep learning models based on other diffusion processes (e.g., Schrödinger-bridge models) for meteorological applications.
@@ Acknowledgements
Appendix A Schrödinger‑bridge model (SM)
@@
Appendix B Diffusion model (DM)
@@
The forward SDE (8) maps the residual $\bm{x}_{\mathrm{HR}}-\bm{x}_{\mathrm{LR}}$ to noise, whereas the reverse SDE (9) maps noise to the residual. The reverse SDE is analytically derived from the forward SDE and includes the score function $\nabla_{\bm{y}_{t}}\ln p_{t}(\bm{y}_{t}|\bm{x}_{\mathrm{LR}},\bm{\xi})$ [3, 14]. If the score function is known, HR samples can be generated from noise by solving the reverse SDE. Specifically, $\bm{x}_{\mathrm{LR}}$ is added to $\bm{y}_{0}$, obtained by solving the reverse SDE, to yield the HR sample. Theoretically, this HR sample follows the conditional distribution $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$.
@@
where $\mu_{t}$ and $\sigma_{t}$ are obtained explicitly from $\lambda_{t}$ [e.g., 12], and $\bm{\eta}\in\mathbb{R}^{n}$ represents Gaussian noise ($\bm{\eta}\sim{\cal N}(0,I_{n})$). The loss function $\mathcal{L}_{\mathrm{DM}}$ for denoising score matching is given by [15]
@@
Appendix C Detailed Comparison of SM and DM
The SM and DM generate HR samples via their respective SDEs, but there are three key differences. First, the SM transforms the point mass $\delta(\bm{x}-\bm{x}_{\mathrm{LR}})$ into $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ [7], whereas the DM transforms ${\cal N}(0,I_{n})$ into the same conditional distribution [40]. Figure 1 in the main text schematically illustrates these transformations. Compared with Gaussian noise lacking spatial structure, $\bm{x}_{\mathrm{LR}}$ is expected to have spatial structure closer to that of HR samples. Owing to this difference in initial conditions, the SM is expected to generate data more efficiently [e.g., 22].
@@ Appendix C Detailed Comparison of SM and DM
The third difference lies in the reference process. SB problems derive optimal transformations in the sense of being closest to reference processes [24]. Both the DM and SM describe the reference processes as SDEs with linear drifts, where the final distributions converge to ${\cal N}(0,I_{n})$ [8, 7]. However, the SM explicitly incorporates $\bm{x}_{\mathrm{LR}}$ in the reference SDE, making the SB problem conditional on $\bm{x}_{\mathrm{LR}}$. In contrast, the DM does not use such conditioning. Conditional methods are particularly effective for large data, such as images [22, 7, 48], and allow more efficient learning and inference compared with unconditional SB problems [8, 6].

### v2

@@ 4.2 Variance for ensemble SR inference
Figure 5 shows scatter plots of RMSE versus spread and rank histograms. Both the DM and SM exhibit spread–skill ratios less than 1 and U-shaped rank histograms, indicating ensemble underdispersion. Such underdispersion is common in deep learning models for meteorological problems [e.g., 27], motivating the development of models with larger spread. Compared with the DM, the SM is better calibrated: its point cloud lies closer to the diagonal, with a spread–skill ratio (Spread/RMSE) of 0.658, larger than 0.649 for the DM. Rank histograms are also flatter for the SM; the Jensen–Shannon distance [10] from a perfectly flat uniform distribution is 0.186 for the SM, smaller than 0.238 for the DM.
Figure 5: (Left) Scatter plots of ensemble spread versus ensemble mean RMSE; (Right) rank histograms. (Top) The DM with $N_{T}=50$; (Bottom) the SM with $N_{T}=10$. The number of ensemble members was 64. In the left column, mean spread–skill ratios (Spread/RMSE) are shown in the figure titles, and in the right column, Jensen–Shannon distances [JS distances; 10] from a uniform distribution are shown. The closer the spread–skill ratio is to 1 and the closer the Jensen–Shannon distance is to 0, the more appropriate the ensemble variance.
To assess significance, we used the results from U-Nets initialized with different random weights and computed the means and standard deviations of the spread–skill ratio and the Jensen–Shannon distance (Table 2). The spread–skill ratio for the SM is closer to 1 than that for the DM, and the Jensen–Shannon distance for the SM is closer to 0. We also computed another common diagnostic, the continuous ranked probability score [CRPS; 14, 51]. The CRPS is smaller for the SM, consistent with the other diagnostics (Table 2). Thus, ensemble inference with the SM yields more appropriate statistics, exhibiting milder underdispersion than the DM.
Table 2: Spread–skill ratios (Spread/RMSE), Jensen–Shannon distances, and continuous ranked probability scores (CRPS) for the DM and SM. These quantities were computed by performing 64-member inference for each ground truth and then averaging spatially and temporally over the test data (540 sets). Standard deviations were obtained as in Table 1.
Spread–Skill Ratio Jensen–Shannon Distance CRPS [K]
DM (64 Members, $N_{T}=50$) 0.641 $\pm$ 0.024 0.234 $\pm$ 0.025 0.151 $\pm$ 0.004
SM (64 Members, $N_{T}=10$) 0.657 $\pm$ 0.013 0.189 $\pm$ 0.015 0.140 $\pm$ 0.003
@@ 4.3 Total inference time
Finally, we discuss the total inference time for 60-min predictions. We first report the wall-clock times for physics-based MSSG simulations, measured on the Earth Simulator at the Japan Agency for Marine-Earth Science and Technology (JAMSTEC), equipped with AMD EPYC 7742 CPUs. On average, HR simulations required 206 min with 256 CPU cores, whereas LR simulations required 6.19 min with 40 CPU cores [53]. We then measured the average SM inference time on a local workstation equipped with an NVIDIA L40S GPU. For 60-min LR data (60 sets), single-member inference took 7.29 s. Thus, the hybrid method combining LR simulation and SM inference completed 60-min predictions of 2-m temperature in 6.31 min, reducing HR computation time to 3.06% (a 32.6-fold speedup). This factor is comparable to values reported for recent surrogate models in urban airflow simulations [39, 33].
@@ 5 Conclusions
We applied a Schrödinger-bridge model [SM; 7] to super-resolve urban 2-m temperature and showed that the SM achieves accuracy comparable to that of a diffusion model [DM; 17], while requiring only about one-fifth the computational cost. We further showed that SM ensembles yield larger spreads and more appropriate statistics than DM ensembles.
Previous studies demonstrate that DMs can handle other variables, such as precipitation and wind velocity [e.g., 15, 27], with variable-specific preprocessing where needed [e.g., log transformation for precipitation; 22]. These findings suggest that Schrödinger-bridge techniques may be extended to other meteorological variables under similar considerations. The present study did not incorporate acceleration methods in order to focus on the original formulations of the SM and DM. Recent results [e.g., 50] suggest that SMs may remain more efficient than DMs even when both are accelerated. Verifying the impact of such acceleration in meteorological applications is an important direction for future work.
@@ Acknowledgements
Appendix A List of abbreviations
Table S1 summarizes the abbreviations used in the manuscript, along with their full terms, listed in alphabetical order.
Table 3: List of abbreviations.
Abbreviation Full Term
CPU Central Processing Unit
CRPS Continuous Ranked Probability Score
DDIM Denoising Diffusion Implicit Model
DM Diffusion Model
GPU Graphics Processing Unit
HR High Resolution
JS Jensen–Shannon
LR Low Resolution
MSE Mean Squared Error
MSSG Multi-Scale Simulator for the Geoenvironment
RMSE Root Mean Square Error
SB Schrödinger Bridge
SDE Stochastic Differential Equation
SI Supplementary Information
SM Schrödinger-bridge Model
SR Super-Resolution
SSIM Structural Similarity Index Measure
[…]
Appendix B Schrödinger-bridge model (SM)
@@
Appendix C Diffusion model (DM)
@@
The forward SDE (8) maps the residual $\bm{x}_{\mathrm{HR}}-\bm{x}_{\mathrm{LR}}$ to noise, whereas the reverse SDE (9) maps noise to the residual. The reverse SDE is analytically derived from the forward SDE and includes the score function $\nabla_{\bm{y}_{t}}\ln p_{t}(\bm{y}_{t}|\bm{x}_{\mathrm{LR}},\bm{\xi})$ [3, 16]. If the score function is known, HR samples can be generated from noise by solving the reverse SDE. Theoretically, this HR sample follows the conditional distribution $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$.
@@
where $\mu_{t}$ and $\sigma_{t}$ are obtained explicitly from $\lambda_{t}$ [e.g., 13], and $\bm{\eta}\in\mathbb{R}^{n}$ represents Gaussian noise ($\bm{\eta}\sim{\cal N}(0,I_{n})$). The loss function $\mathcal{L}_{\mathrm{DM}}$ for denoising score matching is given by [17]
@@
Appendix D Detailed Comparison of SM and DM
The SM and DM generate HR samples via their respective SDEs, but there are three key differences. First, the SM transforms the point mass $\delta(\bm{x}-\bm{x}_{\mathrm{LR}})$ into $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ [7], whereas the DM transforms ${\cal N}(0,I_{n})$ into the same conditional distribution [42]. Figure 1 in the main text schematically illustrates these transformations. Compared with Gaussian noise lacking spatial structure, $\bm{x}_{\mathrm{LR}}$ is expected to have spatial structure closer to that of HR samples. Owing to this difference in initial conditions, the SM is expected to generate data more efficiently [e.g., 24].
@@ Appendix D Detailed Comparison of SM and DM
The third difference lies in the reference process. SB problems derive optimal transformations in the sense of being closest to reference processes [26]. Both the DM and SM describe the reference processes as SDEs with linear drifts, where the final distributions converge to ${\cal N}(0,I_{n})$ [9, 7]. However, the SM explicitly incorporates $\bm{x}_{\mathrm{LR}}$ in the reference SDE, making the SB problem conditional on $\bm{x}_{\mathrm{LR}}$. In contrast, the DM does not use such conditioning. Conditional methods are particularly effective for large data, such as images [24, 7, 50], and allow more efficient learning and inference compared with unconditional SB problems [9, 6]. Finally, we note that the conditional reference process used here follows the specific construction of [7] and represents only one possible choice. There is currently no general guideline for selecting reference processes, and alternative constructions [e.g., 24] may lead to different properties or efficiencies.